OpenStudy (anonymous):

I got Lot Of Maths Tricks.. That U Don't Have.. Feeling This Is Fake.. It worths Millions Of Dollars.. Here Is One Example...

OpenStudy (anonymous):

13\[13\times12 = \] Do u want to take the answer in just 2 seconds..?? \[13\times12 .. \] Make a Circle around 2 Now Add first part of Question.. \[13+2 = 15\] Add Now U Have To Take The Multiply of \[3\times2=6\] Finally U know The Answer Now its = 156 Follow The Steps For To Get The Answer of \[13\] There Are So Many Tricks .. This is A Simple Trick and This Is A Sense :) If u Like To Join Me.. www.facebook.com/RAthish.rulz is open For u..

OpenStudy (anonymous):

uhh

OpenStudy (anonymous):

Or you would already know it if you weren't a 12 year old who didn't know his multiplication table...

OpenStudy (anonymous):

or you could have done \[13\times10=130\]\[13\times2=26\]\[130+26=156\]

hero (hero):

Or this: \((10 + 3)(10 + 2) = 100 + 20 + 30 + 6 = 156\)

OpenStudy (kainui):

Maybe even like this: 13*12=12*12+12=144+12=156

OpenStudy (anonymous):

I Cant Upload All Tricks But If U Join Me U Can Learn it :) rulzrathish@gmail.com is my email <3 #only #maths #lovers